# First-order CP phase transition in two-flavor QCD at $\theta=\pi$ under electromagnetic scale anomaly via a Nambu-Jona-Lasinio description

### v1

@@ IV Summary and discussion
In summary, we have discussed the thermal CP phase transition in QCD at $\theta=\pi$ under a weak magnetic field background with $eB\lesssim m_{\rm dyn}^{2}$, where the electromagnetic scale anomaly plays a significant role. We have worked on a two-flavor PNJL model at $\theta=\pi$ in the MFA, including the electromagnetic-scale anomaly term (Eqs.(14) with (16)). The thermal CP phase transition has been observed to be of first order, in sharp contrast to the two-flavor (P)NJL without the electromagnetic scale anomaly (Figs. 1 and 2).
The electromagnetic-scale anomaly effect is dominated by the thermomagnetic term (the second term in Eq.(16)), which significantly creates the potential barrier at high temperatures around the origin of the CP-order parameter ($P$) direction (Fig. 3). This barrier generation is induced by non-standard potential or nonperturbative form $\sim|eB|^{3}\frac{|P|}{P^{2}+m_{0}^{2}}$ (Eq.(24)), in contrast to the CP-invariant chiral phase transition at $\theta=0$ (Eq.(27)), although both forms are reduced to be identical each other in the chiral limit $m_{0}\to 0$. This implies a significant amplification of the $U(1)$ axial violation due to the electromagnetic scale anomaly at high temperatures when $\theta=\pi$.
Since being insensitive to the electromagnetic scale anomaly, the phase transition along the Polyakov loop direction is not much different from the case without the electromagnetic scale anomaly, therefore goes like smooth crossover with the pseudocriticality. Thus, the presence of the electromagnetic scale anomaly supports the generation of the CP-broken deconfinement(-like) domain $T^{\rm(CP)}_{c}>T^{\rm(dec)}_{\rm pc}$, which becomes wider as $eB$ increases.
@@
An extended Columbia plot at $\theta=\pi$ with varying quark masses under a weak magnetic field background can also straightforwardly be investigated in a way similar to the work in Kawaguchi:2021nsa for the chiral phase transition. In the present analysis, we have also found that the potential barrier generation thermomagneticly induced from the electromagnetic scale-anomaly seems to persist as long as the up and down quark masses are in a range $1010^{-5}<m_{0}/m_{\rm phys}<$ 10, where $m_{\rm phys}$ denotes the value at the physical point. In the case below the lower bound, the CP-even and -odd chiral-order parameters $S$ and $P$ behave like almost identical $(V_{\rm eff}^{\rm(Tad)}\propto 1/|\alpha|)$ in the tadpole term contribution (see Eqs.(27) and (24)), hence the intrinsic barrier structure is gone in the $P$-direction. When the quark mass exceeds the upper bound, no spike structure or nontrivial contribution in the $P$-direction will be left. The parameter space in the two-flavor (P)NJL model outside the first-order regime above would all be the second order regime, unless another explicit CP violation other than what is induced from nonzero $\theta$ is introduced. The case would become nontrivial when the strange quark is incorporated into the model, because of the intrinsic potential barrier structure at $T=0$ associated with the cubic potential term arising from the $U(1)$ axial anomaly. Detailed study on this issue deserves another publication.

### v3

@@ IV Summary and discussion
In summary, we have discussed the thermal CP phase transition in QCD at $\theta=\pi$ under a weak magnetic field background with $eB\lesssim{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Lambda_{\rm QCD}^{2}}$, where the electromagnetic scale anomaly plays a significant role. We have worked on a two-flavor NJL model at $\theta=\pi$ in the MFA, including the electromagnetic-scale anomaly term (Eqs.(12) with (14)). The thermal CP phase transition has been observed to be of first order, in sharp contrast to the two-flavor (P)NJL without the electromagnetic scale anomaly (Figs. 1 and 2).
The electromagnetic-scale anomaly effect is dominated by the thermomagnetic term (the second term in Eq.(14)), which significantly creates the potential barrier at high temperatures around the origin of the CP-order parameter ($P$) direction (Fig. 3). This barrier generation is induced by non-standard potential or nonperturbative form $\sim|eB|^{3}\frac{|P|}{P^{2}+m_{0}^{2}}$ (Eq.(22)), in contrast to the CP-invariant chiral phase transition at $\theta=0$ (Eq.(25)), although both forms are reduced to be identical to each other in the chiral limit $m_{0}\to 0$. This implies a significant amplification of the $U(1)$ axial violation due to the electromagnetic scale anomaly at high temperatures when $\theta=\pi$.
@@
An extended Columbia plot at $\theta=\pi$ with varying quark masses under a weak magnetic field background can also straightforwardly be investigated in a way similar to the work in Kawaguchi:2021nsa for the chiral phase transition. In the present analysis, we have also found that the potential barrier generation thermomagneticly induced from the electromagnetic scale-anomaly seems to persist as long as the up and down quark masses are in a range $1010^{-5}<m_{0}/m_{\rm phys}<$ 10, where $m_{\rm phys}$ denotes the value at the physical point. In the case below the lower bound, the CP-even and -odd chiral-order parameters $S$ and $P$ behave like almost identical $(V_{\rm eff}^{\rm(Tad)}\propto 1/|\alpha|)$ in the tadpole term contribution (see Eqs.(25) and (22)), hence the intrinsic barrier structure is gone in the $P$-direction. When the quark mass exceeds the upper bound, no spike structure or nontrivial contribution in the $P$-direction will be left. The parameter space in the two-flavor NJL model outside the first-order regime above would all be the second order regime, unless another explicit CP violation other than what is induced from nonzero $\theta$ is introduced.
Inclusion of the Polyakov loop variable into the NJL model could give a nontrivial result on the deconfinement phase transition at $\theta=\pi$: the tadpole potential term induced from the electromagnetic scale anomaly would get the Polyakov loop dependence via the Fermi-Dirac distribution function part in $F$ in the thermomagnetic term (see Eq.(16)). Note that the precise form of the Fermi-Dirac distribution function part in $F$ will not substantially be sensitive to the generation of the barrier around the criticality, which acts like almost constant (see Eqs. 21 and 22.), so that the first order nature will still survive with the Polyakov loop extension. Consequently, the thermal evolution of the Polyakov loop variable, would be nontrivially affected by the barrier generation mechanism along the CP order parameter, due to the modification of the peak structure as demonstrated in Eqs.(19) - (22). The deconfinement phase transition might thus become first-order like. Actually, this extension involves a theoretically nontrivial issue on the correlation between the scale anomaly and confinement, and would provide more phenomenological consequences other than the CP phase transition. This issue is thus to be left to be pursued for another publication.
When the strange quark is incorporated into the model, the intrinsic potential barrier structure at $T=0$ associated with the cubic potential term arising from the $U(1)$ axial anomaly. However, the electromagnetic scale-anomaly induced barrier is created near the origin ($P\sim 0$), while the cubic term becomes significant for larger $P$. Therefore, the emergence of the former barrier structure is expected to persist even in the three-flavor case. Detailed study on this issue deserves another publication.